### v1

# LightMem: Lightweight and Efficient Memory-Augmented Generation

Abstract
Despite their remarkable capabilities, Large Language Models (LLMs) struggle to effectively leverage historical interaction information in dynamic and complex environments. Memory systems enable LLMs to move beyond stateless interactions by introducing persistent information storage, retrieval, and utilization mechanisms. However, existing memory systems often introduce substantial time and computational overhead. To this end, we introduce a new memory system called LightMem, which strikes a balance between the performance and efficiency of memory systems. Inspired by the Atkinson–Shiffrin model of human memory, LightMem organizes memory into three complementary stages. First, cognition-inspired sensory memory rapidly filters irrelevant information through lightweight compression and groups information according to their topics. Next, topic-aware short-term memory consolidates these topic-based groups, organizing and summarizing content for more structured access. Finally, long-term memory with sleep-time update employs an offline procedure that decouples consolidation from online inference. Experiments on LongMemEval with GPT and Qwen backbones show that LightMem outperforms strong baselines in accuracy (up to 10.9% gains) while reducing token usage by up to 117×, API calls by up to 159×, and runtime by over 12×. The code is available at https://github.com/zjunlp/LightMem.
1 Introduction
Memory is fundamental to intelligent agent, enabling the assimilation of prior experiences, contextual cues, and task-specific knowledge that underpin robust reasoning and decision-making  (Wang et al., 2024; Behrouz et al., 2024; Du et al., 2025; Zhang et al., 2024). While Large Language Models (LLMs) (DeepSeek-AI et al., 2025; Achiam et al., 2023) demonstrate remarkable capabilities across a wide range of tasks, they exhibit significant limitations when engaged in long-context or multi-turn interaction scenarios due to fixed context windows and the “lost in the middle” problem (Liu et al., 2024). Memory systems are pivotal for overcoming these limitations, as they allow LLMs to maintain a persistent state across extended interactions. Recent works (Li et al., 2025b; Yang et al., 2024; Chhikara et al., 2025; Kang et al., 2025) address this challenge by building explicit external memory through sequential summarization and long term storage, enabling models to retain and retrieve relevant information over long horizons.
Note that a typical LLM memory system processes raw interaction data into manageable chunks, such as turn- or session-level in dialogue scenarios (Xu et al., 2025; Li et al., 2025a), organizes them into long-term memory (e.g., databases or knowledge graphs) by indexing them into memory units, and continuously updates by adding new information and discarding outdated or conflicting content (Zhong et al., 2024). This enables retrieval of relevant memories, improving coherence, and personalization in long-context, multi-turn scenarios.
Challenges. Despite these advances, as shown in Figure 1, contemporary memory systems still suffer from significant inefficiencies and consistency issues. First, in long interactions (e.g., dialogue scenarios), both user inputs and model responses often contain substantial redundant information (Maharana et al., 2024; Wu et al., 2025). Such information is typically irrelevant to downstream tasks or subsequent memory construction, and in some cases, may even negatively affect the model’s in-context learning capability (Liu et al., 2023; Pan et al., 2025). However, current mainstream memory-related studies generally process the raw information directly without any filtering or refinement, leading to high overhead from noisy or irrelevant data. This inflates token consumption without proportional gains in reasoning quality or coherence. Second, memory construction typically treats each turn in isolation or relies on rigid context-window boundaries, failing to model semantic connections across different turns (Tan et al., 2025). As a result, during subsequent memory item construction, the backbone LLM may generate inaccurate or incomplete item representations due to overly entangled topics or semantics, leading to the loss of crucial contextual details. Third, memory updates and forgetting are usually performed directly during inference and task execution. This tight coupling introduces long test-time latency in long-horizon tasks and prevents deeper, reflective processing of past experiences.
In contrast, human memory offers a compelling model of efficiency and adaptability (Liu et al., 2025a). It operates through a hierarchical architecture: sensory memory pre-filters incoming stimuli before they reach short-term memory, which actively integrates, manipulates, and reasons over task-relevant content. Over time, salient information is selectively consolidated into long-term memory processing in sleep time. This dynamic, multi-stage system balances retention, compression, and retrieval, striking an optimal trade-off between performance and resource usage.
Figure 1: Comparison of previous works and LightMem.
Building Lightweight Memory. Inspired by the efficiency and structure of human memory, we introduce LightMem, a lightweight memory architecture designed to minimize redundancy while preserving performance. In particular, LightMem emulates human memory through three key components: (1) A pre-compression sensory memory module that filters redundant or low-value tokens from raw input and buffers the distilled content for downstream processing. This initial filtering step reduces noise before information enters the memory pipeline. (2) A topic-aware short-term memory that leverages semantic and topical similarity to dynamically group related utterances into coherent segments. By adaptively determining segment boundaries based on content instead of fixed window sizes, this module produces more concentrated and meaningful memory units. This not only reduces the frequency of memory construction but also enables more precise and efficient retrieval during inference. (3) A sleep-time update mechanism for long-term memory maintenance. New memory entries are initially stored with timestamps to support immediate (“soft”) updates for real-time responsiveness. Later, during designated offline periods (i.e., “sleep”), the system reorganizes, de-duplicates, and abstracts these entries, resolving inconsistencies and strengthening cross-knowledge connections. Crucially, this decouples expensive memory maintenance from real-time inference, enabling reflective, high-fidelity updates without introducing latency. By systematically filtering, organizing, and consolidating relevant information, LightMem substantially reduces computational overhead and API costs while sustaining accurate, coherent reasoning over extended interactions. We detail each component in §3.
[…]
2 Preliminary
2.1 Conventional Memory Systems for LLMs
Although implementation details and order vary across methods, LLM memory systems can often be formalized as a three-stage model for LLMs: (I) Raw data $D$ are first processed at a chosen level of granularity, $D^{(g)}=f_{\text{gran}}(D;g),g\in\{\text{turn},\text{session},\text{topic}\}$, so that information can be structured to reflect different contextual scopes. (II) The segmented data $D^{(g)}$ are subsequently summarized or extracted to generate memory units, $U=f_{\text{sum}}(D^{(g)})$, which are then stored and organized within structural backends such as vector databases or knowledge graphs, enabling long-term retention. (III) Many systems incorporate an updating mechanism to mitigate issues such as context conflicts or outdated information, $M^{\prime}=f_{\text{update}}(M,R;U)$, where $M$ denotes the existing memory bank, $R$ represents the newly generated memory units, and $U$ specifies the update or forgetting policy.
2.2 Atkinson–Shiffrin Human Memory Model
Figure 2: Human memory system.
Following the Atkinson–Shiffrin human memory model (Atkinson & Shiffrin, 1968), raw environmental information in human brain is first briefly retained in sensory memory as shown in Figure 2, which enables rapid pre-attentive feature extraction and filtering, effectively serving as a form of pre-compression. The processed output can then enter short-term memory (STM), where information and interaction sequences are preserved for tens of seconds to minutes, supporting secondary filtering and more deliberate processing. In contrast, long-term memory (LTM) provides durable storage and undergoes continuous reorganization through updating, abstraction, and forgetting. Importantly, Rasch & Born (2013) highlight that sleep plays a critical role in this reorganization, as oscillatory activity during sleep facilitates the integration and consolidation of memory systems.
2.3 Limitations of Existing LLM Memory Systems
Compared to human memory, current LLM memory systems are burdened by high maintenance costs, mainly due to three limitations: 1) Redundant Sensory Memory. In current systems, $f_{\text{sum}}()$ and $f_{\text{gran}}(;g=\text{topic})$ are typically executed by calling stronger LLMs. Feeding raw data $D$ directly wastes resources and even weakens in-context learning due to redundancy. A key challenge is to design lightweight mechanisms that pre-compress inputs and apply pre-attention strategies to capture semantic units at different granularities efficiently. 2) Balancing Effectiveness and Efficiency in STM. As shown in Figure 1, when input granularity is fixed, $D^{(g)}$ must pass through the entire pipeline. Excessively fine granularity increases latency and underutilizes STM capacity, whereas overly coarse granularity without semantic constraints or grouping may cause mixed or entangled semantics and topics, leading to inaccurate memory construction and loss of fine-grained details in subsequent processes. This calls for strategies that better balance effectiveness and efficiency in STM. 3) Inefficient LTM Updating. Current $f_{\text{update}}()$ mechanisms face two main issues: (i) enforcing strict real-time updates at test time incurs significant latency, whereas STM can provide short-term context without immediate LTM updates; (ii) memory banks are updated sequentially due to ordering constraints (read-after-write/write-after-read), rather than being triggered dynamically. These limitations raise a research question: Can we design LLM memory that is both efficient and lightweight, inspired by human memory mechanisms?
Figure 3: The LightMem architecture. Our LightMem consists of three modules: a) implements an efficient Sensory Memory Module that selectively preserves salient information from raw input, b) realizes a concise STM Module for transient information processing, and c) provides an LTM module designed to minimize retrieval latency.
3 lightmem architecture
Analogous to the human memory, we design LightMem as shown in Figure 3, which consists of three light modules: Light1 implements an efficient Sensory Memory Module that selectively preserves salient information from raw input, Light2 realizes a concise STM Module for transient information processing, and Light3 provides an LTM module designed to minimize retrieval latency.
3.1 Light1: Cognitive-inspired sensory memory
In long horizon interaction scenarios, such as user–assistant dialogues, a large portion of the information is redundant. Therefore, we design a Pre-Compressing Submodule to eliminate redundant tokens, followed by the Topic Segmentation Submodule that forms semantic topic-based segments for following faster and more accurate memory construction.
Pre-Compressing Submodule. This module leverages a compression model $\theta$ to eliminate redundant tokens, tailored for compatibility with the downstream memory construction phase:
$\hat{\mathbf{x}}=\left\{x_{i}\in\mathbf{x}\;\middle|\;P(\text{retain }x_{i}\mid\mathbf{x};\theta)>\tau\right\},\tau=\text{Percentile}\left(\{x_{j}\},r\right),$
[…]
where the subscript $1$ denotes the “retain” class. Tokens with probabilities above a dynamic threshold are included in the compressed sequence. In addition, LightMem can also employ more general generative LLM as the pre-compression model. We further implement a token filtering mechanism based on the cross-entropy between the model’s predicted distribution and the true token labels:
[…]
where $q(x_{i})$ denotes the true token label distribution. Tokens with higher conditional entropy under a given context are more uncertain and less predictable, indicating greater informational uniqueness and a more critical role in semantic expression, such distinctive tokens are essential for subsequent memory construction and are therefore retained.
Topic Segmentation Submodule. Existing works indicate that topic-granular input facilitates improved performance in memory systems (Pan et al., 2025; Tan et al., 2025). As shown in Figure 3, LightMem maintains a sensory memory buffer to temporarily store information after pre-compression. When the accumulated information reaches the buffer’s maximum capacity, a hybrid topic segmentation operation based on attention and similarity is triggered. We use the compression model $\theta$ and an embedding model to compute attention matrices and semantic similarities, respectively. We define the final segmentation boundaries as the intersection of attention-based boundaries $\mathcal{B}_{1}$ and similarity-based boundaries $\mathcal{B}_{2}$:
$\mathcal{B}_{1}$ $=\left\{k\mid M_{k,k-1}>M_{k-1,k-2},\ M_{k,k-1}>M_{k+1,k},\ 1<k<n\right\},$
$\mathcal{B}_{2}=\Bigl\{\,k\;\big|\;\mathrm{sim}(s_{k-1},s_{k})<\tau,1\leq k<n\Bigr\},\quad\mathcal{B}=\mathcal{B}_{1}\cap\mathcal{B}_{2}.$
Specifically, dialogue scenarios possess natural semantic units, namely the conversational turn. We construct a turn-level attention matrix $M\in\mathbb{R}^{n\times n}$. $\mathcal{B}_{1}$ are identified as local maxima in the sequence $\{M_{k,k-1}\}$, i.e., the sub-diagonal elements of $M$ corresponding to attention between consecutive sentences. The detailed process of $\mathcal{B}_{1}$ and illustrative cases are provided in Appendix B.1. To mitigate attention sinks and dilution in attention-based methods, we compute semantic similarity between adjacent turns near each candidate boundary in $\mathcal{B}_{1}$. Boundaries with similarity below threshold $\tau$ form set $\mathcal{B}_{2}$, which helps determine the final topic boundaries $\mathcal{B}$.
3.2 Light2: Topic-aware short-term memory
After obtaining individual topic segments, forming an index structure of {topic, message turns}, where message turns = $user_{i}$, $model_{i}$}. These are first placed into the STM buffer. When the token count in the buffer reaches a preset threshold, we invoke LLM $f_{\text{sum}}$ to generate concise summaries of every structure. The final index structure stored in LTM is {topic, $sum_{i}$, $user_{i}$, $model_{i}$}}.
$\mathrm{sum}_{i}=f_{\mathrm{sum}}\!\left(S_{i}\right),\quad S_{i}\subseteq\{\,\mathrm{user}_{i},\mathrm{model}_{i}\,\},\;S_{i}\neq\varnothing,$
$\mathrm{Entry}_{i}=\left\{\,\mathrm{topic},\;\mathbf{e}_{i}:=\operatorname{embedding}(\mathrm{sum}_{i}),\;\mathrm{user}_{i},\;\mathrm{model}_{i}\,\right\},$
where $\mathrm{Entry}_{i}$ denotes the memory entry to be stored in LTM. Compared with inputting at the granularity of a single turn or session, directly feeding multiple sessions can reduce subsequent API calls but often introduces inaccurate memory entries due to excessive topic mixing, leading to performance degradation. In contrast, topic-constrained input granularity minimizes API calls to the greatest extent while preserving summarization accuracy and maintaining stable system performance.
3.3 Light3: Long-term Memory with Sleep-time Update
Soft Updating at Test Time. At test time, when memory entries arrive, LightMem directly inserts them into LTM with soft updates, thereby decoupling the update process from online inference. Due to real-time updates being converted to direct insertions, interaction latency is significantly reduced. After all entries are inserted or when an update trigger arrives, we compute an update queue for every entry in LTM.
$\mathcal{Q}(e_{i})=\operatorname{Top}_{k}\Big\{(e_{j},\mathrm{sim}(v_{i},v_{j}))\;\mid\;t_{j}\geq t_{i},j\neq i\Big\}_{:n},$
where $e_{i}$ denotes the $i$-th memory entry with embedding $v_{i}$ and timestamp $t_{i}$, $\mathrm{sim}(\cdot,\cdot)$ is the similarity function, and $\operatorname{Top}_{k}\{\cdot\}_{:n}$ indicates selecting the top-$k$ most similar candidates, with the update queue ${Q}(e_{i})$ length fixed at $n$. Consistent with existing work, we select the top-$k$ existing memory entries with the highest semantic similarity as potential update sources. On this basis, we further impose the constraint that only entries with later timestamps are allowed to update earlier ones ($t_{j}\geq t_{i}$), which is consistent with realistic temporal dynamics. Here, $\mathcal{Q}(e_{i})$ denotes the queue of other entries that may update $e_{i}$. Since this process involves only similarity retrieval, it is fast and lightweight, and can be executed offline in parallel with online inference.
Offline Parallel Update. LightMem does not simply transfer online update latency to offline phases, it substantially reduces the overall update latency. The online update mechanism in existing memory frameworks enforces sequential updates, leading to a total latency that accumulates with each update. As shown in Figure 3, in LightMem, each memory entry maintains a global update queue, with each queue corresponding to a distinct $f_{\text{update}}$ operation. Since the update targets are independent across queues, updates can be executed in parallel, thereby greatly reducing the total latency.
Table 1: Effectiveness and efficiency comparison. The token usage is in thousands. – indicates no value for the metric. Bold denotes the best result, underline the second-best. $r$ denotes the compression rate. thth denotes the capacity threshold of the STM buffer, measured in tokens. Each pair of $r$ and thth corresponds to two rows: one for online soft update and one for offline update. OP-update denotes the offline parallel update process of LightMem.
Method ACC (%) Summary Tokens (k) Update Tokens (k) Total (k) Calls Runtime (s)
In Out In Out
[Uncaptioned image] GPT-4o-mini
FullText 56.80 105.07
NaiveRAG 61.00 867.38
LangMem 37.20 982.68 119.48 1,102.16 520.62 2,293.70
A-MEM 62.60 214.66 42.82 1,157.52 190.81 1,605.81 986.55 5,132.06
MemoryOS 44.80 2,302.35 304.18 350.02 35.19 2,991.75 2,938.41 8,030.04
Mem0 53.61 424.13 17.76 560.17 150.56 1,152.62 811.57 4,248.49
LightMem
rr=0.5, thth=256 64.29 20.80 10.01 30.81 25.67 302.69
(OP-update) 64.69 44.46 2.56 47.02 70.23 342.63
rr=0.6, thth=256 67.78 24.58 10.53 35.11 30.47 329.61
(OP-update) 65.39 53.98 3.18 57.16 85.07 411.56
rr=0.7, thth=512 68.64 18.88 9.37 28.25 18.43 283.76
(OP-update) 67.07 79.38 4.06 83.44 125.47 496.03
[Uncaptioned image] Qwen3-30B-A3B-Instruct-2507
FullText 54.80 105.07
NaiveRAG 60.80 659.09
LangMem 50.80 1,311.96 118.06 1,430.02 495.12 3,237.16
A-MEM 65.20 219.21 66.98 1,260.54 318.20 1,864.93 989.30 5,367.51
MemoryOS 49.60 2,101.54 510.88 305.12 27.43 2,944.97 2,922.28 8,721.78
Mem0 39.51 424.20 15.34 411.50 111.35 1001.90 722.76 2,239.94
LightMem
rr=0.4, thth=768 61.95 9.01 16.14 25.15 16.54 357.13
(OP-update) 62.34 111.13 7.88 119.01 176.02 1036.47
rr=0.6, thth=768 70.20 13.19 19.21 32.40 19.97 417.13
(OP-update) 65.14 97.11 5.92 103.03 152.93 1023.56
rr=0.8, thth=1024 68.69 14.82 18.49 33.31 9.43 355.71
(OP-update) 67.34 106.91 6.20 113.11 168.37 1026.90
4 experiments
4.1 experimental setup
Experimental Details. (1) Our experiments adopt a realistic Incremental Dialogue Turn Feeding setting, where dialogue turns are processed sequentially as they arrive. This mirrors practical applications where the full dialogue history is unavailable until the conversation concludes (Hu et al., 2025). (2) For considerations of both efficiency and effectiveness, we employ LLMLingua-2 as our pre-compressor throughout all subsequent experiments. (3) The attention scores for topic segmentation are also obtained using LLMLingua-2; therefore, the size of the sensory memory buffer matches the model’s context window length, which is 512 tokens.
Dataset & Baseline Methods. We use the well-known dataset LongMemEval (Wu et al., 2025) to evaluate memory ability. Specifically, we adopt LongMemEval-S, which consists of 500 dialogue histories, each containing an average of 50 sessions and 110k tokens. Under the evaluation protocol, we choose it because of its balance between dialogue length and computational cost feasibility. We compare LightMem against several representative baselines of conversational memory modeling. ① Full Text, ② Naive RAG, ③ LangMem (LangChain, 2025), ④ A-MEM (Xu et al., 2025), ⑤ MemoryOS (Kang et al., 2025), ⑥ Mem0 (Chhikara et al., 2025). In addition, all methods use GPT-4o-mini and Qwen3-30B-A3B-Instruct-2507 as the LLM backbones. Details on dataset, baselines, and experimental settings are provided in the Appendix C.
Metrics. We evaluate these methods using both effectiveness and efficiency metrics. For effectiveness, we report Accuracy (ACC), defined as the proportion of correctly answered questions. Following prior work, evaluation is conducted with GPT-4o-mini as an LLM judge, guided by a detailed evaluation prompt (see Appendix D.1). For efficiency, we focus on tracking the computational costs of LLM invocations, all averaged across the entire dataset. We define two key operations of memory management: the Summary process as memory distillation and summarization operations, and the Update process as memory consolidation or deletion operations. For both processes, we report the token consumption from LLM calls, including input tokens, output tokens, and total token usage (in thousands). Additionally, we track API Calls counting the total number of LLM invocations, and Runtime recording the execution time in seconds. Our focus is on the overhead of the aforementioned processes. The costs associated with subsequent retrieval and QA are not specifically analyzed, due to their minimal and comparable token/time overhead.
Table 2: The impact of LightMem compression ratio $r$ and STM buffer threshold $th$ is reported here. Due to space limitations, we only present a subset of representative results of the online soft update results, with more results provided in the Appendix 4.
Model thth $r$ ACC Input (k) Output (k) Total (k) Calls Time
GPT
256 0.5 64.29 20.80 10.01 30.81 25.67 302.69
256 0.6 67.68 24.58 10.53 35.11 30.47 329.61
256 0.7 65.68 27.66 9.97 37.63 34.26 403.59
512 0.6 63.74 16.23 9.45 25.68 15.63 266.98
512 0.7 68.64 18.88 9.37 28.25 18.43 283.76
512 0.8 66.67 21.55 8.59 30.14 21.11 268.97
1024 0.6 59.68 10.34 7.68 18.20 7.69 177.45
1024 0.7 64.68 12.93 6.90 19.83 8.25 209.12
1024 0.8 64.35 14.86 6.28 21.14 9.43 216.08
Qwen
512 0.4 58.57 11.03 17.00 28.03 10.11 421.74
512 0.6 66.57 16.22 19.50 35.72 15.40 471.09
512 0.8 67.37 21.35 19.36 40.71 20.98 461.02
768 0.4 61.95 9.01 16.14 25.15 6.54 357.13
768 0.6 73.20 13.19 19.21 32.40 9.97 417.13
768 0.8 64.95 16.94 19.06 36.00 13.09 420.14
1024 0.4 53.91 8.02 15.44 23.46 4.83 300.56
1024 0.6 65.67 11.50 18.21 29.71 7.18 396.35
1024 0.8 68.69 14.82 18.49 33.31 9.43 355.71
4.2 main results
As shown in Table 1, LightMem demonstrates superior effectiveness and efficiency across three parameter settings on both GPT and Qwen backbones. We evaluated LightMem in two stages: one before the sleep time update, i.e., at the end of the online soft update phase when the memory construction is completed; and the other after the offline update.
[…]
Across all three parameter settings, LightMem achieves superior performance on nearly all metrics and both LLM backbones, demonstrating its robustness and high flexibility.
Figure 4: Analysis and Ablation Study of Key Modules. Fig.(a) depicts the QA accuracy when using prompts compressed at different ratios ($r$) as in-contexts to query the LLM directly. Fig.(b) compares the accuracy of different topic segmentation methods under these varying compression ratios. Fig.(c1) and Fig.(c2) present the ablation study for the topic segmentation module, evaluating its impact on both performance and efficiency for the GPT and Qwen models.
4.3 Analysis of Pre-Compressing Submodule
Performance and Overhead. LightMem uses an additional model (Pan et al., 2024b; Xia et al., 2025) for pre-compression. We evaluate its performance by randomly sampling 1/5 of LongMemEval and compressing it at ratios shown in Figure 4(a), then prompting LLMs for in-context QA. When compression ratio $r$ ranges from 50%–80%, compressed and uncompressed performance are comparable, demonstrating LLMs can effectively understand compressed content and validating LightMem’s approach. The submodule is highly efficient, consuming under 2GB of GPU memory with negligible impact on overall runtime.
Impact of $r$ on Performance. As shown in Tables 2 and  4, The optimal $r$ for ACC is dependent on the STM buffer threshold $th$. For smaller thresholds ($th\in\{0,256\}$), an $r$ of 0.6 achieves the highest ACC. In contrast, for larger thresholds ($th\in\{512,1024\}$), a higher retention rate of $r=0.7$ performs best. This suggests greater buffer capacity enables effective use of richer, less-compressed information, leveraging LLMs’ advanced long-context processing to mitigate the “lost in the middle” phenomenon. On average, the optimal $r$ for ACC is 0.6, reflecting a trade-off between information compression rate and the quantity of information in the STM buffer. In terms of efficiency, a lower $r$ generally leads to higher efficiency, as it triggers the buffer threshold less frequently under the same $th$, resulting in fewer API calls and lower token consumption.
4.4 Analysis of Topic Segmentation Submodule
Segmentation Accuracy. To validate the accuracy of our proposed hybrid topic segmentation method, we compare it with segmentation using only a single granularity: attention-only-based and similarity-only-based segmentation. Since the construction process of the LongMemEval indicates that different sessions naturally serve as topic boundaries, we directly use them as ground-truth labels. The final accuracy is calculated as the number of correctly identified segmentation points divided by the total number of labels. The results in Figure 4(b) validate the effectiveness of our method: it achieves higher accuracy than both individual segmentation methods across all compression ratios, with an absolute accuracy exceeding 80%.
Ablation Study. As shown in Figure 4(c), removing the topic segmentation submodule slightly improves efficiency but significantly harms accuracy, causing a 6.3% drop for GPT and 5.4% for Qwen. This indicates that the submodule effectively enables models to perceive semantic units in the input, facilitating subsequent memory unit generation.
4.5 Analysis of the STM Threshold’s Impact
As illustrated in the Figure 5, the STM buffer threshold ($th$) has a distinct but significant impact on both efficiency and performance metrics. A consistent trend is: as $th$ increases, there is a marked improvement in efficiency. In contrast, the effect on QA accuracy is non-monotonic. The optimal threshold for accuracy varies depending on the model and the compression ratio ($r$), indicating that a larger buffer does not always yield better performance. This highlights a crucial trade-off: while a larger STM threshold is consistently better for reducing computational cost, the ideal setting for maximizing task accuracy requires careful tuning.
4.6 Analysis of Sleep-Time Update
Why Soft Updates Work. A primary challenge in designing memory systems is handling updates. While powerful, LLMs can be unreliable when tasked with complex real-time update operations. For instance, when presented with two related but not contradictory pieces of information, an LLM might incorrectly interpret them as a conflict and delete the older memory entry, leading to irreversible information loss. Instead, the optimal operations might be to merge the information or simply add the new entry. In contrast, LightMem performs only incremental additions through soft updates during test time, which preserves global information and complete semantics.
Case Study: Memory Update Mechanism Comparison History1: {’Monday, 2 PM’: User is planning a trip to Tokyo.} History2: {’Monday, 4 PM’: User asks about trains to Kyoto.} Hard Update: Overwrites memory -> "User plans Kyoto trip" .  Tokyo context lost LightMem Soft Update: Appends info -> "Tokyo trip + Kyoto inquiry" Full context preserved
5 related work
Hard Prompt Compression for LLMs. Hard prompt compression improves LLM efficiency by removing redundant content from prompts (Li et al., 2025c). Methods recently have evolved from using smaller language models (Jiang et al., 2023; Li et al., 2023; Chuang et al., 2024) to query-aware approaches that preserve task-relevant information (Weston & Sukhbaatar, 2023; Creswell et al., 2023; Jiang et al., 2024). Additionally, lightweight bidirectional encoders have demonstrated strong effectiveness and efficiency (Pan et al., 2024a; Liskavets et al., 2025).
Chunking Strategies in RAG Systems. Retrieval-Augmented Generation (RAG) systems rely on chunking extrernal documents into smaller units for retrieval (Lewis et al., 2020; Gao et al., 2023). Existing chunking strategies include rule-based methods creating fixed-size segments (Lewis et al., 2020; Sarthi et al., 2024; Edge et al., 2024; Gutierrez et al., 2024), semantic-based methods grouping content by topic (Qu et al., 2025), and LLM-driven methods leveraging model knowledge for splitting (Pan et al., 2025; Duarte et al., 2024; Zhao et al., 2024; Liu et al., 2025b). However, all of these chunking strategies for RAG systems are tailored to static scenarios, not applicable to dynamic and open-ended environments.
Memory Systems for LLM Agents. Memory systems help LLM agents move beyond stateless interactions to support flexible reasoning and adaptation in complex and changing environments (Liu et al., 2025a; Mei et al., 2025). The earliest and most straightforward approaches store experiences as linear or sequential streams, sometimes enhanced with hierarchical structures (Liang et al., 2023; Park et al., 2023; Packer et al., 2023; Zhong et al., 2024; Salama et al., 2025; Fang et al., 2025). A more structured class of methods represents memories as nodes and their relationships as edges, using trees, graphs, or temporal knowledge structures to support retrieval and update (Rezazadeh et al., 2025; Chhikara et al., 2025; Rasmussen et al., 2025; Xu et al., 2025; Zhang et al., 2025). The latest trend integrates various types of memory, allowing them to interact and synergistically improve overall performance (Kang et al., 2025; Li et al., 2025b; Wang & Chen, 2025; Nan et al., 2025). Overall, existing memory systems for LLM agents have become increasingly complex and capable, leveraging hierarchical, structured, and multi-type memories. However, most focus on maximizing effectiveness, with limited consideration of efficiency. While some recent works (Guo et al., 2024; Zhao et al., 2025; Dong et al., 2025) share a similar motivation with our work, they focus on lightweight adaptations of GraphRAG where the corpus is predefined and static.
6 conclusion and Future Work
In this work, we introduced LightMem, a lightweight and efficient memory framework designed to address the significant overhead of memory systems for LLM agents. Inspired by the multi-stage Atkinson-Shiffrin human memory model, LightMem’s architecture effectively filters, organizes, and consolidates information. Our empirical evaluation demonstrates that this approach maintains strong task performance while sharply reducing computational costs.
Offline Update Acceleration. We plan to enhance the efficiency of LightMem’s update phase by incorporating pre-computed key–value (KV) caches. This optimization allows KV cache pre-computation to be performed offline, significantly accelerating memory consolidation and reducing runtime overhead during interactive sessions.
Knowledge Graph-based Memory. To better address complex multi-hop reasoning challenges, we aim to integrate a lightweight knowledge graph-based memory module. This module will support explicit relational reasoning and structured information retrieval, enabling the model to perform more interpretable and compositional reasoning across interconnected knowledge entities.
Multimodal Memory Extension. We further plan to develop a multimodal memory mechanism that allows LightMem to adapt to multimodal models and scenarios. Such an extension is crucial for embodied agents and real-world applications where visual, auditory, and textual information jointly contribute to memory formation and retrieval.
Parametric–Nonparametric Synergy. We will explore the collaborative mechanisms between parametric and non-parametric memory components. To bridge the gap between the two paradigms, we aim to enable more flexible, synergistic knowledge utilization, combining the efficiency of parametric representations with the interpretability and adaptability of non-parametric storage.
Ethics Statement
LightMem enhances LLM agents by creating an external memory of user interactions. While this improves agent coherence, it introduces critical ethical challenges. Storing dialogue histories poses inherent risks to user privacy, as conversations may contain sensitive data. The memory can also absorb and perpetuate biases or misinformation from user input, potentially leading to bad agent behavior. Therefore, any deployment of this technology must prioritize robust safeguards. We strongly advocate for strict privacy protocols, such as data anonymization and user consent, as well as mechanisms to mitigate the effects of biased or false memories. Responsible development is essential to ensure these memory-augmented systems are used in a safe and trustworthy manner.
Reproducibility Statement
To ensure the reproducibility of this work, we introduce the detailed implementations for LightMem are provided in in Section 3, Appendix B. Additionally, we plan to release our source code in the future to further support reproducibility. These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
Appendix A Usage of LLMs
Throughout the preparation of this manuscript, we used LLMs to assist with improving grammar, clarity, and wording in parts of this work. The use of LLMs was limited to language refinement, with all ideas, analyses, and conclusions solely developed by the authors.
Figure 5: Impact of the STM buffer threshold ($th$) on performance and efficiency across different compression ratios ($r$). Each radar chart represents a specific configuration of a model (GPT-4o-mini or Qwen3) and a fixed compression ratio. The axes measure six key metrics: Accuracy (ACC), token consumption (Input, Output, Total), API Calls, and Runtime. To facilitate comparison, all values are normalized for visualization on the chart.
Appendix B Methodology Details
B.1 Topic Segmentation
In this part, we present the construction of the attention matrix, the underlying rationale for topic segmentation, and representative illustrative cases.
We extract only the user sentences from multi-turn dialogues, as they are generally more concise and the assistant’s responses necessarily remain consistent with the user’s theme within the same turn. Moreover, since the maximum input length of the LLMLingua-2 Pan et al. (2024b) model is 512 tokens, the assistant’s often lengthy sentences cannot be effectively accommodated. Therefore, we sequentially store the user sentences into a buffer and segment them, ensuring that as many sentences as possible are preserved while staying within the token limit. As a practical trick, if a sentence becomes empty after compression, we retain its original uncompressed version; if the token length of a sentence still exceeds the maximum limit, we continue to compress it using the LLMLingua-2 model at a 0.5 compression rate until the token length falls below the threshold. To reduce the effect of attention sinks, we mask out the contributions of the first and last three tokens in each sequence and subsequently normalize the remaining attention values. Attention is derived from the higher layers of LLMLingua-2 (layers 8, 9, 10, and 11). For any two sentences, we first compute token-level pairwise attention and average across tokens to obtain the overall attention of one sentence to the target sentence; we then average across the selected layers to obtain a more robust inter-sentence attention score. For each current sentence, the attention scores directed toward all preceding sentences are normalized within the sentence, yielding the final attention matrix. Residual fragments that remain after segmentation are carried over to the beginning of the next buffer for further processing, and this procedure continues iteratively until the dialogue ends.
Based on the attention pattern, we focus on the sequence formed by each sentence’s attention scores relative to its immediately preceding sentence, which directly reflects the continuity of local semantics. Therefore, we take the attention scores from the outermost layer of the attention map. When the attention score at a given position is higher than both its preceding and following positions, it is regarded as a local peak. If a sentence is identified as a peak, we set a segmentation point immediately before this sentence, making the peak sentence the beginning of a new segment. The rationale is that the peak sentence exhibits consistently low attention to all earlier sentences overall and reflects a clear transition from an old topic to a new one, indicating that the identified sentence marks the initiation of a new topic.
Figure 6: Example of Topic Segment Attention Matrix.
Figure 6 illustrates three representative examples of reliable segmentation under 50% compression rate. In the first attention map, local peaks in the adjacent-sentence attention sequence appear at positions 5, 8, and 11, where the actual segmentation boundaries lie between sentences 4–5 and 11–12. In the second attention map, peaks occur at positions 3, 8, 12, and 14, and the actual boundaries are located between sentences 7–8, 11–12, and 13–14. Overall, our method achieves close alignment with the majority of true boundaries while providing finer-grained segmentation. These examples demonstrate that our segmentation approach enables both fine-grained and reliable detection of topic boundaries, thereby validating its effectiveness.
B.2 Category-wise Accuracy
As summarized in Table 3, retrieval-augmented and memory-centric methods (e.g., A-MEM, Mem0, MemoryOS) generally outperform Full Text on categories that demand information integration or belief revision, such as Temporal, Multi-Session, and Knowledge-Update. In contrast, categories such as Single-User and Single-Assistant, lightweight retrieval like Naive RAG is often competitive and can be the most reliable option, while Single-Preference shows higher variance due to its smaller sample size.
Table 3: Category-wise Accuracy. Accuracy (%) by method across question types. Parentheses indicate category proportion and sample size. For GPT, LightMem is configured with parameters $r=0.7$ and $\text{th}=512$; for Qwen, LightMem is configured with $r=0.4$ and $\text{th}=768$.
Method
Temporal
( n=133n{=}133)
Multi-Session
(n=133n{=}133)
Knowledge-Update
(n=78n{=}78)
Single-User
(n=70n{=}70)
Single-Assistant
(n=56n{=}56)
Single-Preference
(n=30n{=}30)
[Uncaptioned image] GPT-4o-mini
Full Text 31.58 45.45 76.92 87.14 89.29 36.67
Naive RAG 39.85 48.48 67.95 90.00 98.21 53.33
LangMem 15.79 20.30 66.67 60.00 46.43 60.00
A-MEM 47.36 48.87 64.11 92.86 96.43 46.67
MemoryOS 32.33 31.06 48.72 80.00 64.29 30.00
Mem0 40.15 46.21 70.12 81.43 41.07 60.00
LightMem 67.18 71.74 83.12 87.14 32.14 68.18
[Uncaptioned image] Qwen3-30B-A3B-Instruct-2507
Full Text 33.08 35.61 76.92 82.86 87.50 50.00
Naive RAG 36.84 47.73 65.38 91.43 98.21 70.00
LangMem 37.60 38.35 67.95 78.57 42.86 70.00
A-MEM 51.88 51.12 76.93 90.00 96.43 40.00
MemoryOS 28.57 36.84 61.54 72.86 92.86 33.33
Mem0 41.94 28.13 28.57 55.32 26.09 81.82
LightMem 54.20 51.91 66.67 80.00 31.25 80.00
B.3 Detailed Parameter Analysis
As Table 4 shows, we report the numerical results of the effects of LightMem parameters (compression ratio $r$ and STM threshold $th$).
Table 4: The impact of LightMem’s compression ratio ($r$) and STM buffer threshold (thth).
Model thth $r$ ACC Input (k) Output (k) Total (k) Calls Time
GPT-4o-mini
0 0.4 58.04 27.70 8.90 36.60 39.91 500.69
256 0.4 57.78 16.64 8.40 25.04 20.25 254.93
512 0.4 55.56 11.05 7.66 18.71 10.13 230.59
768 0.4 49.29 9.05 6.55 15.60 6.57 157.13
1024 0.4 46.87 7.75 5.25 13.00 4.82 118.11
0 0.5 62.89 30.84 9.75 40.59 43.56 550.36
256 0.5 64.29 20.80 10.01 30.81 25.67 302.69
512 0.5 62.44 13.49 8.89 22.38 12.70 250.36
768 0.5 56.12 10.93 7.57 18.50 8.12 203.13
1024 0.5 50.36 8.34 6.97 15.31 6.32 160.35
0 0.6 70.35 33.17 10.20 43.37 45.86 553.07
256 0.6 67.68 24.58 10.53 35.11 30.47 329.61
512 0.6 63.74 16.23 9.45 25.68 15.63 266.98
768 0.6 64.44 13.04 8.10 21.14 9.90 210.05
1024 0.6 59.68 10.34 7.68 18.20 7.69 177.45
0 0.7 62.35 35.36 9.76 45.12 48.08 573.42
256 0.7 65.68 27.66 9.97 37.63 34.26 403.59
512 0.7 68.64 18.88 9.37 28.25 18.43 283.76
1024 0.7 64.68 12.93 6.90 19.83 8.25 209.12
0 0.8 61.52 39.32 9.89 49.21 52.97 622.90
256 0.8 66.37 30.67 9.70 40.37 41.66 489.61
512 0.8 66.67 21.55 8.59 30.14 21.11 268.97
1024 0.8 64.35 14.86 6.28 21.14 9.43 216.08
Qwen3
0 0.4 56.89 28.44 18.30 46.74 41.08 594.94
256 0.4 52.37 16.82 17.63 34.45 20.48 450.98
512 0.4 58.57 11.03 17.00 28.03 10.11 421.74
768 0.4 61.95 9.01 16.14 25.15 6.54 357.13
1024 0.4 53.91 8.02 15.44 23.46 4.83 300.56
0 0.6 69.56 34.90 20.26 55.16 48.63 642.10
256 0.6 65.37 24.78 19.59 44.37 30.66 520.37
512 0.6 66.57 16.22 19.50 35.72 15.40 471.09
768 0.6 73.20 13.19 19.21 32.40 9.97 417.13
1024 0.6 65.67 11.50 18.21 29.71 7.18 396.35
0 0.8 67.68 37.97 20.18 58.15 50.81 759.15
256 0.8 64.52 30.54 19.77 50.31 37.35 550.98
512 0.8 67.37 21.35 19.36 40.71 20.98 461.02
768 0.8 64.95 16.94 19.06 36.00 13.09 420.14
1024 0.8 68.69 14.82 18.49 33.31 9.43 355.71
Appendix C Experiment Details
C.1 Datasets and Baselines
The LongMemEval dataset (Wu et al., 2025) is designed to benchmark long-term interactive memory in conversational agents. It comprises 500 evaluation questions built upon extended user-assistant dialogues. It has two different versions with different lengths: the LongMemEval-S setting contains approximately 115k tokens per problem, while the LongMemEval-M setting extends up to 1.5 million tokens across 500 sessions. In our work, we adopt the LongMemEval-S version due to its balance between dialogue length and computational feasibility. Each evaluation instance consists of a multi-session dialogue history, a question posed at a later timestamp, and the corresponding ground truth answer with supporting evidence spans. The questions are categorized into multiple types: information extraction, multi-session reasoning, knowledge updates, temporal reasoning, and abstention. Overall, the dataset is characterized by extremely long histories, wide temporal spans, and diverse question types, making it a comprehensive benchmark for evaluating conversational agents’ memory capabilities. During the experiments, five samples from this dataset contained corrupted characters, which caused LightMem’s compression model to fail to run properly. Consequently, LightMem directly discarded these five samples when processing the dataset. However, their accuracy results were uniformly treated as false. The indices of these five samples in the dataset are 74, 183, 278, 351, and 380.
We compare our approach against several representative baselines of conversational memory modeling. ① LangMem (LangChain, 2025): The Langchain’s long-term memory module. ② A-MEM (Xu et al., 2025): Constructs a memory-centric knowledge graph, encoding each interaction as a structured memory note and linking these notes via LLM-driven reasoning. ③ MemoryOS (Kang et al., 2025): Organizes conversational memory in an OS-inspired hierarchy, structuring interactions into short-term, mid-term, and long-term layers via paging and heat-based updating. ④ Mem0 (Chhikara et al., 2025): Extracts memories from dialogue turns through a combination of global summaries and recent context, maintaining them via LLM-guided operations.
C.2 Implementation Details
All the experiments are conducted on hardware equipped with 4×NVIDIA RTX 3090 GPUs, dual Intel Xeon Gold 6133 CPUs (40 cores, 80 threads), and 256 GB of RAM.
Appendix D Prompts
D.1 LLM-as-Judge
Standard Tasks (Single-session-user/assistant  Multi-session) I will give you a question, a correct answer, and a response from a model. Please answer yes if the response contains the correct answer. Otherwise, answer no. If the response is equivalent to the correct answer or contains all the intermediate steps to get the correct answer, you should also answer yes. If the response only contains a subset of the information required by the answer, answer no. Question: {question} Correct Answer: {answer} Model Response: {response} Is the model response correct? Answer yes or no only.
Temporal Reasoning Tasks I will give you a question, a correct answer, and a response from a model. Please answer yes if the response contains the correct answer. Otherwise, answer no. If the response is equivalent to the correct answer or contains all the intermediate steps to get the correct answer, you should also answer yes. If the response only contains a subset of the information required by the answer, answer no. In addition, do not penalize off-by-one errors for the number of days. If the question asks for the number of days/weeks/months, etc., and the model makes off-by-one errors (e.g., predicting 19 days when the answer is 18), the model’s response is still correct. Question: {question} Correct Answer: {answer} Model Response: {response} Is the model response correct? Answer yes or no only.
Knowledge Update Tasks I will give you a question, a correct answer, and a response from a model. Please answer yes if the response contains the correct answer. Otherwise, answer no. If the response contains some previous information along with an updated answer, the response should be considered as correct as long as the updated answer is the required answer. Question: {question} Correct Answer: {answer} Model Response: {response} Is the model response correct? Answer yes or no only.
Single-session Preference Tasks I will give you a question, a rubric for desired personalized response, and a response from a model. Please answer yes if the response satisfies the desired response. Otherwise, answer no. The model does not need to reflect all the points in the rubric. The response is correct as long as it recalls and utilizes the user’s personal information correctly. Question: {question} Rubric: {answer} Model Response: {response} Is the model response correct? Answer yes or no only.
Abstention Tasks I will give you an unanswerable question, an explanation, and a response from a model. Please answer yes if the model correctly identifies the question as unanswerable. The model could say that the information is incomplete, or some other information is given but the asked information is not. Question: {question} Explanation: {answer} Model Response: {response} Does the model correctly identify the question as unanswerable? Answer yes or no only.